In trapezoid PQRS, Base PQ is parallel to base RS. Let point X be the intersection of diagonals PR and QS. The area of triangle PQR is 4 and the area of triangle QRX is 4. Find the area of trapezoid PQRS
Found 2 solutions by CPhill, ikleyn:
Answer by CPhill(2103)

This is an interesting problem that utilizes the properties of triangles within a trapezoid.
Here is how to find the area of trapezoid PQRS.
## 📐 Given Information
* **Trapezoid PQRS** with base **PQ** parallel to base **RS** ($PQ \parallel RS$).
* Diagonals **PR** and **QS** intersect at point **X**.
* Area of $\triangle PQR$ is **4**.
* Area of $\triangle QRX$ is **4**.
## 🧐 Analysis and Calculation
### Step 1: Find the Area of $\triangle PQX$
The area of $\triangle PQR$ is the sum of the areas of $\triangle PQX$ and $\triangle QRX$:
$$Area(\triangle PQR) = Area(\triangle PQX) + Area(\triangle QRX)$$
We are given $Area(\triangle PQR) = 4$ and $Area(\triangle QRX) = 4$.
$$4 = Area(\triangle PQX) + 4$$
$$Area(\triangle PQX) = 4 - 4$$
$$Area(\triangle PQX) = 0$$
This result seems unusual for a geometric figure. Let's **re-examine the given information**. If $Area(\triangle PQR) = 4$ and $Area(\triangle QRX) = 4$, and $\triangle QRX$ is a part of $\triangle PQR$, the only way this is possible is if $\triangle PQR$ and $\triangle QRX$ are the **same triangle**, which would mean point **P** coincides with point **X**.
However, if P coincides with X, then the diagonal PR would just be the single point P, which doesn't intersect with QS at a distinct point X, unless Q, P, and R are collinear, which is not the case in a standard trapezoid.
---
### Alternative Interpretation of the Prompt
Given the geometric context, it is highly likely that there is a **typo** in the problem statement and the intended area was for $\triangle **RSX**$ or $\triangle **PSX**$.
Let's assume the given areas are for two triangles that **share the same height** or **base** properties, a common setup in these problems.
A common scenario that makes these problems solvable is when the given areas are for **adjacent** triangles that are **not** subsets of each other, for example, $Area(\triangle PQX)$ and $Area(\triangle QRX)$, or $Area(\triangle PQX)$ and $Area(\triangle PSX)$.
Let's **assume the intended prompt meant:**
$$Area(\triangle PQX) = 4$$
$$Area(\triangle QRX) = 4$$
*(This is a standard setup where $Area(\triangle PQR) = 8$.)*
### Step 2: Use the Equal Area Property
In any trapezoid, the two triangles formed by the legs and the segment of the diagonals have equal area. That is, $Area(\triangle PSX)$ and $Area(\triangle QRX)$ are equal.
$$Area(\triangle PSX) = Area(\triangle QRX)$$
Since we assumed $Area(\triangle QRX) = 4$:
$$Area(\triangle PSX) = 4$$
### Step 3: Find the Ratio of Areas
In $\triangle PQS$, the base $PQ$ is not parallel to $RS$. $\triangle PQX$ and $\triangle PSX$ share the same vertex $P$ and their bases $QX$ and $SX$ lie on the line $QS$. The ratio of their areas is equal to the ratio of their bases:
$$\frac{Area(\triangle PQX)}{Area(\triangle PSX)} = \frac{QX}{SX}$$
$$\frac{4}{4} = \frac{QX}{SX} \implies \frac{QX}{SX} = 1$$
This means $X$ is the midpoint of $QS$.
Now consider $\triangle PQR$ and $\triangle S R X$ which is wrong. Let's consider $\triangle PQX$ and $\triangle RSX$. These two triangles are **similar** because $PQ \parallel RS$, which means $\angle QPX = \angle RSX$ and $\angle PQX = \angle SRX$ (alternate interior angles), and $\angle PXQ = \angle RXS$ (vertical angles).
Since $\triangle PQX \sim \triangle RSX$, the ratio of their corresponding heights is the same as the ratio of their bases:
$$\frac{Height\text{ of } \triangle PQX}{Height\text{ of } \triangle RSX} = \frac{PQ}{RS}$$
The ratio of the areas of two similar triangles is the square of the ratio of their corresponding sides:
$$\frac{Area(\triangle PQX)}{Area(\triangle RSX)} = \left(\frac{PQ}{RS}\right)^2$$
### Step 4: Find $Area(\triangle RSX)$
Consider $\triangle PQR$ and $\triangle S R Q$ which is wrong.
Consider $\triangle PQX$ and $\triangle RSX$. They are similar.
We know that for triangles sharing a height (like $\triangle PQX$ and $\triangle RSX$ sharing a height from $R$ is wrong), the ratio of their areas is the ratio of their bases.
Let $h_1$ be the height of $\triangle PQX$ from $X$ to $PQ$.
Let $h_2$ be the height of $\triangle RSX$ from $X$ to $RS$.
We use the ratio property from the diagonals intersecting at $X$:
$$\frac{Area(\triangle PQX)}{Area(\triangle RSX)} = \frac{Area(\triangle PRQ) - Area(\triangle PSQ)}{Area(\triangle PSR) - Area(\triangle QSR)} \text{ is wrong}$$
Using the ratio of bases derived from the common heights:
1. In $\triangle PQS$: $\frac{Area(\triangle PQX)}{Area(\triangle PSX)} = \frac{QX}{SX}$
$$\frac{4}{4} = 1 \implies \mathbf{QX = SX}$$
2. In $\triangle QRS$: $\frac{Area(\triangle QRX)}{Area(\triangle RSX)} = \frac{QX}{SX}$
Since $QX = SX$:
$$\frac{Area(\triangle QRX)}{Area(\triangle RSX)} = 1$$
$$Area(\triangle RSX) = Area(\triangle QRX)$$
Since $Area(\triangle QRX) = 4$:
$$\mathbf{Area(\triangle RSX) = 4}$$
### Step 5: Find the Area of the Trapezoid
The area of the trapezoid is the sum of the areas of the four triangles formed by the diagonals:
$$Area(\text{Trapezoid } PQRS) = Area(\triangle PQX) + Area(\triangle QRX) + Area(\triangle PSX) + Area(\triangle RSX)$$
$$Area(\text{Trapezoid } PQRS) = 4 + 4 + 4 + 4$$
$$Area(\text{Trapezoid } PQRS) = 16$$
---
## Conclusion
Assuming the problem intended to give **$Area(\triangle PQX) = 4$** and **$Area(\triangle QRX) = 4$** (to lead to a standard and solvable geometry problem), the area of the trapezoid is:
$$Area(\text{Trapezoid } PQRS) = **16**$$
If we strictly adhere to the initial prompt where $Area(\triangle PQR) = 4$ and $Area(\triangle QRX) = 4$, the problem is either **impossible** or implies a **degenerate** trapezoid where $Area(\triangle PQX)=0$, which is not meaningful for this type of question. The interpretation above is the standard solution for a very similar problem.
Would you like me to find similar geometry problems or check the math for a different set of assumed areas?

Answer by ikleyn(53250)

.
In trapezoid PQRS, Base PQ is parallel to base RS. Let point X be the intersection of diagonals PR and QS.
The area of triangle PQR is 4 and the area of triangle QRX is 4. Find the area of trapezoid PQRS
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


The post is SELF-CONTRADICTORY and describes a situation which NEVER may happen.


Indeed, triangle QRX is a part of triangle PQR, and triangle QRX is a proper,
i.e., smaller part of triangle PQR.

Due to this reason, the areas of triangles PQR and PRX can not be equal.

This is a contradiction, which kills the problem to the death and ruins it into dust.